OpenStudy (anonymous):

@johnweldon1993 (let me get the pic up) but would I be right about this?

OpenStudy (anonymous):

\[5\sqrt{3}+\sqrt{3} would this be 5\sqrt{6}\]

OpenStudy (johnweldon1993):

No, you wouldn't add the numbers in the square roots...those stay the same and just get left alone \[5 \sqrt{3} + \sqrt{3} = (5 + 1)\sqrt{3}\]

OpenStudy (johnweldon1993):

General rule: \[\large a\sqrt{c} + b\sqrt{c} = (a + b)\sqrt{c}\]

OpenStudy (anonymous):

\[6\sqrt{3}\]

OpenStudy (johnweldon1993):

That would be correct! :)